OpenStudy (anonymous):

Algebra Slope Question Determine which function has the greater rate of change in problems 1−3.

OpenStudy (anonymous):

A The Slopes are Equal B The Graph has a greater slope C The equation has a greater slope D None of the above

OpenStudy (anonymous):

y = 2x + 7

OpenStudy (anonymous):

y = 2x + 7 -- In y = mx + b form, the slope is in the m position...therefore, the slope of this equation is 2. As for the graph...pick any two sets of points and find the slope by using the slope formula : (y2 - y1) / (x2 - x1) I am picking : (1,3)...x1 = 1 and y1 = 3 (-1,-1)...x2 = -1 and y2 = -1 now we sub (-1 - 3) / (-1 - 1) -4/-2 which reduces to 2 so the slope of the points on the graph is also 2. SLOPES ARE EQUAL any questions ?